Reports 1-1 of 1 Clear search Modify search
MIF (General)
takafumi.ushiba - 15:59 Tuesday 28 January 2025 (32475) Print this report
Comment to Investigation of PRMI_3F lock (32467)

[Tanaka, Ushiba]

We checked HWP angle and found that the current angle is 164 (fig1) while the value in the medm is 168.
So, we changed medm values to 164 so that the value matches to the real hardware.

Though HWP angle has a mismatch, it doesn't explain the power reduction on REFL because we need to rotate it by 8 degrees to adjust the power on REFL PD to the old value.

Note:

We found that PDA2 screw was loose and easy to move, so we fixed them and realign the beam to PDA2. 

Images attached to this comment
Search Help
×

Warning

×